OpenStudy (anonymous):

A bag contains 7 blue marbles, 8 green marbles, 15 red marbles, and 10 yellow marbles. Find the probability of randomly drawing without replacement, a blue marble and then a red. a. 7/160 or about 0.0438 b. 21/320 or about 0.0656 c. 7/104 or about 0.0673 d. 15/66 or about 0.2273

OpenStudy (anonymous):

i think its b) 21/320

OpenStudy (turingtest):

I actually get 21/319

OpenStudy (turingtest):

21/312 I mean :/

OpenStudy (anonymous):

are you sure? when you divide 1600 by 5 you get 320 right?

OpenStudy (turingtest):

it says "without replacement" so the way I see it (and I can be wrong) the first drawing is 7/40 and the second is 15/39 (because on the second drawing there are only 39 marbles in the bag) either way that is very close to b)

OpenStudy (anonymous):

hmm youre right but its still close to b

OpenStudy (turingtest):

oh yeah, that's a choice I forgot how to simplify :) no, I'm pretty sure I'm right

OpenStudy (turingtest):

you can simplify my answer, and it does not simplify to b)

OpenStudy (anonymous):

right but i think you should still choose be since it is closest to 21/319

OpenStudy (turingtest):

no, that was a typo my answer is \[\frac7{40}\times\frac{15}{39}\]

OpenStudy (turingtest):

simplify...

OpenStudy (turingtest):

does that reasoning make sense to you?

OpenStudy (anonymous):

yes so if you simplfy you get 105/1560 then if you divide by 5 you get 21/319

OpenStudy (turingtest):

no, check your arithmetic the whole point is that there are fewer marbles in the bag on the second drawing

OpenStudy (anonymous):

hmm but isnt the question asking the probablity of picking a red then blue with out replacement??

OpenStudy (turingtest):

right, no replacement means 40 marbles in the bag the first time, and 39 in the bag the second make sense?

OpenStudy (turingtest):

if there was replacement that would mean we put the first blue marble back, and there would be 40 marbles again in the bag for the second drawing without putting that blue marble we pulled out back the bag is 1 marble short; 39

OpenStudy (anonymous):

right

OpenStudy (turingtest):

so now that you see where the fractions come from, just compute\[\frac7{40}\times\frac{15}{39}\]and be careful with your simplification